OpenStudy (anonymous):

Find the dimensions of x and y that maximizes the area given that the perimeter is =100

OpenStudy (anonymous):



OpenStudy (anonymous):

help

OpenStudy (anonymous):

I will add more info

OpenStudy (anonymous):

Area= xy+piy^2/4

OpenStudy (anonymous):

Perimeter=piy+2x

OpenStudy (anonymous):

Satellite can u plz help me?

OpenStudy (anonymous):

yeah we can do this

OpenStudy (anonymous):

YAY

OpenStudy (anonymous):

U r the best

OpenStudy (anonymous):

you have the perimeter as \[P=100=2x+\pi y\] right?

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

solve for y to get \[y=\frac{100-2x}{\pi}\]

OpenStudy (anonymous):

area is \[xy+\pi(\frac{y}{2})^2\] yes?

OpenStudy (anonymous):

yes i just got frozen sorry

OpenStudy (anonymous):

just checking because i am going slow and don't want to screw up the algebra